1,w,wsquare are the cube roots of unity (1+w) (1+wsquare) (1+wseven) (1+w8)=1 proove the sum

Grade:12

1 Answers

Vijay Mukati
7 years ago
Dear Student,

w^7 = w and w^8 = w^2 and thus given question can be simplified to:
[(1 + w)(1 + w^2)]]^2 = {1 + w + w^2 + w^3}^2 = 1
Hence proved.

Thanks.